# Exploiting the diversity of modeling methods to probe systematic biases in strong lensing analyses

### v1

@@ Acknowledgements.
This work originated in the Lensing Odyssey 2021 workshop, and so we would like to acknowledge the organizers and attendees for the fruitful discussions. AG acknowledges the Swiss National Science Foundation (SNSF) for supporting this work. GV and QM were both supported by the generosity of Eric and Wendy Schmidt by recommendation of the Schmidt Futures program. QM gratefully acknowledges a grant of computer time from ACCESS allocation TG-AST130007. This research has made use of SciPy (Virtanen et al. 2020), NumPy (Oliphant 2006; Van Der Walt et al. 2011), Matplotlib (Hunter 2007), Astropy (Astropy Collaboration et al. 2013; Price-Whelan et al. 2018), Jupyter (Kluyver et al. 2016) and GetDist (Lewis 2019).
@@ Appendix A Realistic source surface brightness in mock data
Figure 13: Old simulated lensing data based on another source galaxy. For that mock the source image background was insufficiently subtracted, which left a faint imprints in the data after being lensed. While unnoticeable to the eye in the data (left panel), the square boundaries of the source are clearly noticeable in the input source (middle panel) and even on the reconstructed source (right panel, using a correlated field model as an example). These artificial boundaries led to biases in lens potential parameters and source size. Note that a logarithmic scale is used in all panels.
@@ Appendix B Technical modeling details
Table 3: Computation times for the models, including best-fit optimization and posterior distribution sampling, obtained with the six different modeling methods. Note that some of the final models were obtained by combining several model instances, so we quote the runtime for a single instance (the most expensive) in addition to the total runtime (assuming single instances were not performed in parallel). For additional details and software package references, see Sect. 3 and Appendix B.
@@
This section gives technical details on the specific choices made by modelers to analyse the simulated data. We complement it with Table 3 that give the approximate computation time necessary to obtain the various models used in this work.
@@ B.2 Adaptive+Matérn
Finally, we further restrict the parameters to 10 per cent of their full range and run a final model with the same regularization but even higher resolution, i.e. shooting back every pixel in the data image to create the adaptive grid. We combine the last two models, which differ only in the resolution of the reconstructed source (although the first served to initialize the second in order to save on computations), by merging their posterior samples in such a way that the probability mass of each one is the same independently of its actual size.
@@ B.3 Cluster+Exp and Cluster+Exp+Lumweight
The Cluster+Exp+Lumweight model uses all of the methods described above, but in addition it uses a luminosity-weighted regularization as described in Sect. 3.3. Thus we include the additional parameter $\rho$ (Eq. 13), which controls the steepness of the luminosity weighting, as an additional nonlinear parameter to be varied.
@@ B.5 Correlated Field
Similar to the strategy used with the multi-scale source model, we first model the imaging data with a single Sérsic source profile, which we then replace with the correlated field model defined in Eq. 15. For our baseline model, we set the shape of the Gaussian excitation field to $10090^{2}=8^{\prime}$ 100 pixels. The field power spectrum is modeled as power law parametrized with an amplitude and a slope. These two parameters are themselves sampled from a log-normal distribution, described in turn by a mean and scale parameters. The additive offset in real space is modeled by a single scalar (initialized to the mean flux values from the Sérsic model), whereas variations of this offset are sampled from a log-normal distribution with additional mean and scale parameters. Lens potential model parameters (PEMD and external shear) are sampled from Gaussian distribution, for which we check that the prior widths are large enough. We optimize the full model—lens mass parameters and source field parameters—using the set of minimizers and samplers implemented in NIFTy. More specifically, we converge to the maximum a posteriori solution and estimate the joint posterior distribution using metric-Gaussian variational inference (MGVI, Knollmüller & Enßlin 2019).
We run several variations of the above fiducial models. In particular, we increase the field resolution to $400120^{2}=14^{\prime}$ 400 pixels, alter the sampler random seeds, initialize the model with a worst lens model (obtained from a different Sérsic source model), and use geometric variational inference (geoVI, Frank et al. 2021) instead of MGVI. All these model variations result in almost identical posterior distributions for lens mass parameters and consistent source models. Nevertheless, we conservatively marginalize over these models with equal weights. For the point-estimate parameters, we set those to the mean values of the VI samples of the model with the most resolved source (i.e., $120^{2}$ pixels), although the fiducial model is virtually indistinguishable.

### v2

@@ Acknowledgements.
The authors thank the anonymous referee for reviewing the original manuscript and providing useful comments. This work originated in the Lensing Odyssey 2021 workshop [^11] https://lensingodyssey.com, and so we would like to acknowledge the organizers and attendees for the fruitful discussions. AG acknowledges the Swiss National Science Foundation (SNSF) for supporting this work. This work was also supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (COSMICLENS: grant agreement No 787886). GV and QM were both supported by the generosity of Eric and Wendy Schmidt by recommendation of the Schmidt Futures program. QM gratefully acknowledges a grant of computer time from ACCESS allocation TG-AST130007. This research has made use of SciPy (Virtanen et al. 2020), NumPy (Oliphant 2006; Van Der Walt et al. 2011), Matplotlib (Hunter 2007), Astropy (Astropy Collaboration et al. 2013; Price-Whelan et al. 2018), Jupyter (Kluyver et al. 2016) and GetDist (Lewis 2019).
@@ Appendix A Realistic source surface brightness in mock data
Figure 13: Old simulated lensing data based on another source galaxy. For that mock the source image background was insufficiently subtracted, which left faint imprints in the data after being lensed. While unnoticeable to the eye in the data (left panel), the square boundaries of the source are clearly noticeable in the input source (middle panel) and even on the reconstructed source (right panel, using a correlated field model as an example). These artificial boundaries led to biases in lens potential parameters and source size. Note that a logarithmic scale is used in all panels.
@@ Appendix B Technical modeling details
Table 3: Computation times for the six modeling methods, including best-fit optimization and posterior distribution sampling.
@@
121212 Some of the final models were obtained by combining several model instances, so we quote the runtime for a single instance (the most expensive) in addition to the total runtime (assuming single instances were not performed in parallel). For additional details and software package references, see Sect. 3 and Appendix B.
This section gives technical details on the specific choices made by modelers to analyze the simulated data. We complement it with Table 3 which gives the approximate computation time necessary to obtain the various models used in this work.
@@ B.2 Adaptive+Matérn
Finally, we further restrict the parameters to 10 per cent of their full range and run a final model with the same regularization but even higher resolution; that is, shooting back every pixel in the data image to create the adaptive grid. We combine the last two models, which differ only in the resolution of the reconstructed source (although the first served to initialize the second in order to save on computations), by merging their posterior samples in such a way that the probability mass of each one is the same independently of its actual size.
@@ B.3 Cluster+Exp and Cluster+Exp+Lumweight
The Cluster+Exp+Lumweight model uses all of the methods described above, but in addition it uses a luminosity-weighted regularization, as is described in Sect. 3.3. Thus we include the additional parameter $\rho$ (Eq. 13), which controls the steepness of the luminosity weighting, as an additional nonlinear parameter to be varied.
@@ B.5 Correlated Field
Similar to the strategy used with the multi-scale source model, we first model the imaging data with a single Sérsic source profile, which we then replace with the correlated field model defined in Eq. 15. For our baseline model, we set the shape of the Gaussian excitation field to $10090^{2}=8^{\prime}$ 100 pixels. The field power spectrum is modeled as power law parametrized with an amplitude and a slope. These two parameters are themselves sampled from a log-normal distribution, described in turn by a mean and scale parameters. The additive offset in real space is modeled by a single scalar (initialized to the mean flux values from the Sérsic model), whereas variations in this offset are sampled from a log-normal distribution with additional mean and scale parameters. Lens potential model parameters (PEMD and external shear) are sampled from Gaussian distribution, for which we check that the prior widths are large enough. We optimize the full model—lens mass parameters and source field parameters—using the set of minimizers and samplers implemented in NIFTy. More specifically, we converge to the maximum a posteriori solution and estimate the joint posterior distribution using metric-Gaussian variational inference (MGVI, Knollmüller & Enßlin 2019).
We ran several variations of the above fiducial models. In particular, we increased the field resolution to $400120^{2}=14^{\prime}$ 400 pixels, alter the sampler random seeds, initialized the model with a worst lens model (obtained from a different Sérsic source model), and used geometric variational inference (geoVI, Frank et al. 2021) instead of MGVI. All these model variations result in almost identical posterior distributions for lens mass parameters and consistent source models. Nevertheless, we conservatively marginalize over these models with equal weights. For the point-estimate parameters, we set those to the mean values of the VI samples of the model with the most resolved source (i.e., $120^{2}$ pixels), although the fiducial model is virtually indistinguishable.